# Blockchain with proof of quantum work

### v2

@@ D.1 Blockchain Efficiency and Transaction Delay
We define blockchain efficiency as the expected number of blocks in the strongest chain (path) divided by the total number of blocks broadcast, in the limit of many blocks. For viable parameterizations this concentrates at a finite value as the number of block broadcasts becomes large. For purposes of Figs. 5 and 16 we estimate this value once after every block broadcast for blockchains of between $512$ and $1024$ broadcasts (mining events). We discard the values for the first 50% of the chain evolution and average the remaining result. The resulting value is shown as a mean and standard error with respect to 16 independent chains.
We can establish two such independent strongest paths, using two stakeholders who have not mined any blocks, and define their pair delay, $D_{pair}$, as the number of blocks in the first path absent from the second path (i.e. the depth at which there is disagreement in the strongest chain). If two users agree on the strongest chain this delay is 00, and is otherwise positive. Stakeholders can be miners: the pair delay describes the distribution of mining power relative to a typical user. If pair-delays are with high probability less than $\mathcal{D}$, then a stakeholder can have corresponding confidence that blocks of depth at least $\mathcal{D}$ on the strongest chain are immutable, since mining power is concentrated on branches compatible with the block validity. We call $\mathcal{D}$ the transaction delay.
For figure 16 we measure 1 pair-delay per chain following each mining event on chains of length from 512 up to 1024. We determine the empirical distribution of $D_{pair}$, discarding statistics from the first half of each chain. We threshold the distribution to determine a transaction delay with 95% confidence, presenting the mean and variance with respect to 16 blockchains as indicated in captions. Although pair-delays measured this way are correlated, for each data point we have at least thousands of effectively independent samples for each data point, so that a 95% quantile is established with reasonable confidence [60].
@@ D.2 Mining Acceleration for Blockchain Analysis
Each mining event requires $O(2^{\mathcal{N}_{\rm zeros}})$ experiments. For purposes of blockchain structural analysis we show methods whereby only $O(1)$ experiments are required per mining event. This allows us to present, with practical compute resources, results on stability for large $\mathcal{N}_{\rm zeros}$, limited primarily by the number of validation experiments rather than the exponentially growing mining timescale.
@@
We first make the approximation in all experiments of perfect network synchronization, discounting the rare event of two miners simultaneously completing a proof-of-work. This approximation is reasonable if the mining rate is small enough.
We employ a hashing method where the overall sign on any hyperplane is uniformly distributed in $\{-1,1\}$. For this reason the probability a nonce describes a broadcastable (valid) block in case (a) is $1/2^{-\mathcal{N}_{\rm zeros}}$; mining is successful if and only if the hyperplane signs $H_{\alpha}=\mathrm{sign}(W_{\alpha})$. We can therefore emulate properly distributed statistics of a fair mining event by first running an experiment to produce $W_{\alpha}$, and then conditionally sampling the signs on hyperplanes to allow mining.
For a confidence-based Chainwork, the probability of successful mining is a function of more than the signs on the hyperplanes. To accelerate mining events in this case, we perform a large number of experiments to determine many copies of $W_{\alpha}$ for $\alpha=1$ to $\mathcal{N}_{\rm zeros}$. In each case, we exhaustively enumerate all hyperplane signs consistent with a mining event. We thereby enumerate a subset of all mining events. These mining events are correlated (in some cases differing only by the signs on hyperplanes), but this is judged to be a weak correlation. Most often only one value for the signs is supported, and events that allow many sign values are rare for the parameter ranges presented. The rate of mining can be inferred, and the mining subset can be used in a randomized order to model the mining.
@@ D.3 Bootstrapping Witness Methodology
For the case that Chainwork is defined as +/-1 for matching (unmatching) sequences, knowing the distribution of the $\mathcal{N}_{\rm zeros}$ bits for each device used in mining and validation is sufficient to evaluate the distribution of Chainwork and thereby the evolution of the blockchain. For the case of confidence-based Chainkwork, it is sufficient to know the distribution of $\{W_{\alpha}\}$.
@@
For a training set of 5 experiments and $1024$ random projections, the distribution of bits (thence validation rates), and distribution of $W_{\alpha}$ (thence confidence in validation) can be established as shown in Figs. 13 and 14. We can fit the witness distributions by Gaussians, and thresholded bit outcomes by Bernouilli distributions. For purposes of accelerating analysis we can then resample these distributions uniformly at random to model a larger set of experiments without use of the QPU experiment each time.
@@
We employ this procedure in Figs. 5 and 16 with results consistent with full blockchain experiments involving finite-number of miners, such as those in Fig. 10. We consider 20480 combinations of model (spin-glass realization) and hyperplanes, for each of the 4 experimental QPUs. Using 32 programmings, we determine the distribution (Gaussian or Bernouilli for witnesses and bits respectively). We can then use these parameterized distributions to generate reasonable witness (and thresholded witness) distributions. For each mining attempt, we uniformly randomly sample a combination of model and hyperplanes - and then for any mining or validation event we choose uniformly at random a QPU and sample from the corresponding distribution.
@@
We can further save on computation costs by only mining (validating) as necessary for purposes of constructing the blockchain. We need only understand a miner’s verification pattern if they are responsible for broadcasting the next block. We therefore construct on-the-fly a miner’s Chainwork and establish their strongest chain at the time of mining. With this approximation we can work in the informative limit of well-distributed compute power (many miners), where each miner mines at most one block in any blockchain. The number of validation events required to realize a chain with $L$ blocks broadcast in total is $L(L-1)/2$.
@@
Qualitative justification for this approach relies on the observation that a spin-glass ensemble operated in our experimental regime is self-averaging. The distributional properties of such witnesses fluctuate little between typical spin-glass disorder realizations. Variation in witnesses values between QPUs (and embeddings) appears well described by Gaussian distributions, and since hyperplanes are nearly orthogonal the correlations in their fluctuations are weak. We plot in Figs. 5 both results obtained by bootstrapping and those obtained by other experiments (either with fixed number of miners, or many miners) with good agreement, validating the approximation. Full simulations (every verification and mining event uses a QPU) are used for all other figures.

### v3

@@ D.1 Blockchain Efficiency and Transaction Delay
We define blockchain efficiency as the expected number of blocks in the strongest chain (path) divided by the total number of blocks broadcast, in the limit of many blocks. For viable parameterizations this concentrates at a finite value as the number of block broadcasts becomes large. A per-blockchain realization value can be estimated by averaging values obtained between broadcast (mining event) $512$ and $1024$. Figs. 5 and 16 show the mean and standard deviation with respect to $16$ independent blockchains.
We can establish two such independent strongest paths, using two stakeholders who have not mined any blocks, and define their pair delay, $D_{pair}$, as the number of blocks in the first path absent from the second path (i.e. the depth at which there is disagreement in the strongest chain). If two users agree on the strongest chain this delay is 0, and is otherwise positive. Stakeholders can be miners: the pair delay describes the distribution of mining power relative to a typical user. If pair-delays are with high probability less than $\mathcal{D}$, then a stakeholder can have corresponding confidence that blocks of depth at least $\mathcal{D}$ on the strongest chain are immutable, since mining power is concentrated on branches compatible with the block validity. We call $\mathcal{D}$ the transaction delay.
For figure 16 we measure 1 pair-delay per chain following each mining event on chains of length from 512 up to 1024. We determine the empirical distribution of $D_{pair}$, discarding statistics from the first half of each chain. We threshold the distribution to determine a transaction delay with 95% confidence, presenting the mean and variance with respect to 16 blockchains as indicated in captions. Although pair-delays measured this way are correlated, for each data point we have at least thousands of effectively independent samples, so that a 95% quantile is established with reasonable confidence [64].
@@ D.2 Mining Acceleration for Blockchain Analysis
Each mining event requires $O(2^{\mathcal{N}_{\rm zeros}})$ nonce attempts (quantum computations). For purposes of blockchain structural analysis we show methods whereby only $O(1)$ nonces are required per mining event. This allows us to present, with practical compute resources, results on stability for large $\mathcal{N}_{\rm zeros}$, limited primarily by the number of validating computations rather than the exponentially growing mining timescale.
@@
We first make the approximation of perfect network synchronization, discounting the rare event of two miners simultaneously completing a proof-of-work. This approximation is reasonable if the mining rate is small enough.
We employ a hashing method where the overall sign on any hyperplane is uniformly distributed in $\{-1,1\}$. For this reason the probability a nonce describes a broadcastable (valid) block in case (a) is $1/2^{-\mathcal{N}_{\rm zeros}}$; mining is successful if and only if the hyperplane signs $H_{\alpha}=\mathrm{sign}(W_{\alpha})$. We can therefore emulate properly distributed statistics of a fair mining event by first calculating $W_{\alpha}$ for a uniformly sampled random nonce, and then sampling the hyperplane signs conditioned upon the $\mathcal{N}_{zeros}$ requirement.
For a confidence-based Chainwork, the probability of successful mining is a function of more than the signs on the hyperplanes. To accelerate mining events in this case, we compute a large number of quantum hashes to determine many copies of $W_{\alpha}$ for $\alpha=1$ to $\mathcal{N}_{\rm zeros}$. In each case, we exhaustively enumerate all hyperplane signs consistent with a mining event. We thereby enumerate a subset of all mining events. These mining events are correlated (in some cases differing only by the signs on hyperplanes), but this is judged to be a weak correlation. Most often only one value for the signs is supported, and events that allow many sign values are rare for the parameter ranges presented. The rate of mining can be inferred, and the mining subset can be used in a randomized order to model the mining.
@@ D.3 Bootstrapping Witness Methodology
For the case that Chainwork is defined as +/-1 for matching (unmatching) sequences, knowing the distribution of the $\mathcal{N}_{\rm zeros}$ bits for each device used in mining and validation is sufficient to evaluate the distribution of Chainwork and thereby the evolution of the blockchain. For the case of confidence-based Chainwork, it is sufficient to know the distribution of $\{W_{\alpha}\}$.
@@
For a training set of $20$ nonce values (unitary evolution parameterizations) and $1024$ random projections, the distribution of bits (thence validation rates), and distribution of $W_{\alpha}$ (thence confidence in validation) can be established as shown in Figs. 13 and 14. We can fit the witness distributions by Gaussians, and thresholded bit outcomes by Bernoulli distributions. For purposes of accelerating analysis we can then resample these distributions uniformly at random to model a larger set of quantum hashes without use of the QPU experiment each time.
@@
We employ this procedure in Figs. 5 and 16 with results consistent with full blockchain demonstrations involving finite-number of miners, such as those in Fig. 10. We consider 20480 combinations of model (spin-glass realization) and hyperplanes, for each of the 4 generally-available QPUs. Using 32 programmings, we determine the distribution (Gaussian or Bernoulli for witnesses and bits respectively). We can then use these parameterized distributions to generate reasonable witness (and thresholded witness) distributions. For each mining attempt, we uniformly randomly sample a combination of model and hyperplanes, and then for any mining or validation event we choose uniformly at random a QPU and sample from the corresponding distribution.
@@
We can further save on computation costs by only mining (validating) as necessary for purposes of constructing the blockchain. We need only understand a miner’s verification pattern if they are responsible for broadcasting the next block. We therefore construct on-the-fly a miner’s Chainwork and establish their strongest chain at the time of mining. With this approximation we can work in the informative limit of well-distributed compute power (many miners), where each miner mines at most one block in any blockchain. The number of validation events required to realize a chain with $L$ blocks broadcast in total is L(L1)/2L(L-1)/2.
@@
Qualitative justification for this approach relies on the observation that the spin-glass ensemble of unitary evolutions in our demonstration is self-averaging. The distributional properties of such witnesses fluctuate little between typical spin-glass disorder realizations. Variation in witnesses values between QPUs (and embeddings) appears well described by Gaussian distributions, and since hyperplanes are nearly orthogonal the correlations in their fluctuations are weak. We plot in Figs. 5 both results obtained by bootstrapping and those obtained in demonstrations (either with fixed number of miners, or many miners) with good agreement, validating the approximation. Full simulations (every verification and mining event uses a QPU) are used for all other figures.